# MILo: Mesh-In-the-Loop Gaussian Splatting for Detailed and Efficient Surface Reconstruction

### v1

@@ 6. Experiments
No dataset provides ground truth geometry for background objects. To mitigate this issue we propose to measure the visual consistency between the extracted mesh and reference views, which contain background information. We thus propose a novel experimental setup, which we coin Mesh-Based Novel View Synthesis (see Section. 6.3). This evaluation measures the ability of reconstructed meshes to accurately represent full scenes, including background objects, while requiring only a set of ground truth images.
@@
Table 1 provides a comprehensive comparison of the computational resources required by different methods.
@@ 6.2. Surface Reconstruction
We provide results on the DTU dataset, consisting of small, object-centric scenes. Table 3 shows the results on this dataset. Our mesh-in-the-loop optimization maintains competitive performance on DTU, even though it is designed for unbounded scenes.
Table 4. Mesh-Based Novel View synthesis metrics on real-world datasets. We report PSNR, SSIM, and LPIPS metrics, as well as the total number of Gaussians and vertices (in millions). Our method demonstrates superior performance on these challenging unbounded scenes.
@@
Figure 6. Reconstruction $\rightarrow$ Ground truth Distance Evaluation. The T&T (Knapitsch et al., 2017) evaluation toolkit reports cumulative histograms (c) showing the proportion of reconstructed points within a given distance from the ground truth surface. The dashed line marks the threshold used to compute the F1-score. Our method produces significantly more points close to the true surface (within a normalized distance of $1.0$), demonstrating higher reconstruction accuracy compared to prior approaches. Qualitatively, our reconstructions using the differentiable from RaDe-GS
(a) are closer to the ground truth point clouds than both GOF (b) and RaDe-GS.
@@ 6.3. Mesh-Based Novel View Synthesis
Despite extensive research on mesh reconstruction from images, current evaluation benchmarks remain limited. Common datasets—DTU, MipNeRF 360, and Tanks & Temples—all have notable drawbacks: MipNeRF 360 lacks ground-truth geometry; DTU features overly controlled, object-centric scenes; and Tanks & Temples provides sparse ground-truth limited to foreground regions. Moreover, there exists no evaluation protocols to measure the alignment of extracted surfaces with ground truth images.
We raise this is an important issue in the surface reconstruction task. To address this, we introduce Mesh-Based Novel View Synthesis, a novel image-based evaluation method for meshes that relies solely on ground-truth RGB images. The core intuition is that better geometry should yield better mesh-based renderings: for each test view, we render the scene using the mesh and compare the resulting image to the corresponding ground-truth image.
@@
In practice, for each mesh under evaluation, we first train the neural field using only the training views. We adopt TensoRF(Chen et al., 2022) as the backbone representation and optimize it for 5k iterations. We then render test views for evaluation. This metric effectively quantifies the alignment between mesh geometry and the image data.
Table 4 shows how Mesh-Based Novel View Synthesis allows us to compare against other state-of-the-art methods. Our approach outperforms 2DGS (Huang et al., 2024) and RaDe-GS (Zhang et al., 2024) across all metrics. GOF (Yu et al., 2024) produces highly detailed meshes, yielding strong PSNR results, but our method achieves comparable PSNR while outperforming in SSIM and LPIPS. This suggests that our meshes are not only accurate but also exhibit fewer visual artifacts and less noise compared to those from GOF. We include a qualitative comparison in the supplementary material.
@@ 6.4. Novel View Synthesis
Figure 7. Effect of adding Normal supervision. Despite showcasing an apparent decrease in performance in our ablation study (see Supplemental), this figure makes supports our statement that the combination of depth and normal supervision is essential. The resulting mesh when using only $\mathcal{L}_{\text{MD}}$ is significantly noisier than when supervising with $\mathcal{L}_{\text{mesh}}$.
@@ 6.5. Ablation Study
To evaluate the contribution of each component, we conduct ablation studies on the Tanks and Temples dataset (see Tab. 5). Adding depth supervision via the mesh loss ($+\mathcal{L}_{\text{MD}}$) substantially boosts $F_{1}$-score, while further including the normal loss ($+\mathcal{L}_{\text{mesh}}$) slightly reduces it. Still, Fig. 7 shows that combining both is critical for removing noise from meshes.
Introducing the anti-erosion loss ($+\mathcal{L}_{\text{mesh}}+\mathcal{L}_{\text{erosion}}$) improves reconstruction by preserving thin structures and fine details—especially in challenging regions like fences or vegetation—by penalizing surface erosion. Our loss also acts as a mechanism to recover geometry lost due to SDF sign flips yielding a modest $F_{1}$-score gain (Fig. 5).
@@ 6.6. Limitations and Conclusions
Future work could explore novel adaptive sampling strategies for Delaunay sites, and improved initialization techniques. Additionally, extending our approach to dynamic scenes and exploring its applications in real-time rendering and interactive applications would be promising directions for future research. Further exploration of surface-based processing tools that can be integrated into the training pipeline could also yield significant improvements in reconstruction quality and versatility. Finally, our mesh-based novel view synthesis evaluation underscores the lack of standardized quantitative protocols to evaluate the alignment of the reconstructed surface with the training images. This is true on widely used datasets such as DTU, MipNeRF 360, and Tanks & Temples. While our approach offers a step toward addressing this gap, it remains an initial effort; future work could focus on refining the evaluation methodology to better capture reconstruction quality in challenging real-world scenes.
Refer to caption Refer to caption
7. Acknowledgement
Parts of this work were supported by the ERC Consolidator Grant “VEGA” (No. 101087347), the ERC Advanced Grant “explorer” (No. 101097259), the ERC Advanced Grant “NERPHYS” (No. 101141721) as well as the ANR AI Chair AIGRETTE.

### v2

@@ 6. Experiments
No dataset provides ground truth geometry for background objects. To mitigate this issue we use Mesh-Based Novel View Synthesis  (see Section. 6.3) to measure the visual consistency between the extracted mesh and reference views, which contain background information. This evaluation measures the ability of reconstructed meshes to accurately represent full scenes, including background objects, while requiring only a set of ground truth images.
@@
The addition of mesh-based constraints inevitably leads to increased computational load. However, we observed that only a subset of Gaussians is critical for producing an effective tetrahedral structure.
Building on this, inspired by the sampling strategy introduced in Mini-Splatting2 (Fang and Wang, 2024b), we employ it for a distinct purpose: selecting the Gaussians which will spawn the Gaussian pivots that serve as Delaunay vertices. Unlike Mini-Splatting2, our dense model preserves all Gaussians throughout optimization, and the sampling criterion is used exclusively to identify those most suitable for mesh construction rather than to prune the representation.
This targeted selection yields lighter and more accurate meshes, significantly reducing the computational cost of triangulation, Marching Tetrahedra, and mesh rasterization. Table 1 provides a comprehensive comparison of the computational resources required by concurrent methods.
@@ 6.2. Surface Reconstruction
We provide results on the DTU dataset, consisting of small, object-centric scenes. Table 3 shows the results on this dataset. Note that, DTU consists of isolated objects in highly controlled scenes, where standard post-hoc mesh extraction is already effective. MILo is particularly effective for complex full-scene reconstruction. Still, our mesh-in-the-loop optimization maintains competitive performance in terms of Chamfer Distance (CD) on DTU. In line with established practice, we evaluate performance using F-Score on the Tanks and Temples dataset and Chamfer Distance (CD) on DTU.
Table 4. Mesh-Based Novel View synthesis metrics on real-world datasets. We report PSNR, SSIM, and LPIPS metrics, as well as the total number of Gaussians and vertices (in millions). Our method demonstrates superior performance on these challenging unbounded scenes. Ours (base) uses the differentiable rasterizer from RaDe-GS.
@@
Figure 6. Reconstruction $\rightarrow$ Ground truth Distance Evaluation. Qualitatively, our reconstructions using the differentiable rasterizer from RaDe-GS (a) are closer to the ground truth point clouds than both GOF (b) and RaDe-GS. The T&T (Knapitsch et al., 2017) evaluation toolkit reports cumulative histograms (c) showing the proportion of reconstructed points within a given distance from the ground truth surface. The dashed line marks the threshold used to compute the F1-score. Our method produces significantly more points close to the true surface (within a normalized distance of $1.0$), demonstrating higher reconstruction accuracy compared to prior approaches.
@@ 6.3. Mesh-Based Novel View Synthesis
Despite extensive research on mesh reconstruction from images, current evaluation benchmarks remain limited. Common datasets—DTU, MipNeRF 360, and Tanks & Temples—all have notable drawbacks: MipNeRF 360 lacks ground-truth geometry; DTU features overly controlled, object-centric scenes; and Tanks & Temples provides sparse ground-truth limited to foreground regions. Moreover, evaluation protocols to measure the alignment of extracted surfaces with ground truth images remain uncommon in surface reconstruction research.
We highlight this as an important issue in surface reconstruction. Inspired by previous works which evaluate surface-based view synthesis methods (Reiser et al., 2024), we use Mesh-Based Novel View Synthesis as an evaluation method for meshes that relies solely on ground-truth RGB images. The core intuition is that better geometry should yield better mesh-based renderings: for each test view, we render the scene using the mesh and compare the resulting image to the corresponding ground-truth image.
@@
In practice, for each mesh under evaluation, we first train the neural field using only the training views. We adopt TensoRF(Chen et al., 2022) as the backbone representation and optimize it for 5k iterations. We then render test views for evaluation. This metric effectively quantifies the alignment between mesh geometry and the image data. Note that, this evaluation protocol assumes dense testing views, as achieving good mesh-based rendering performance on a sparse test set does not necessarily translate into high quality of the underlying geometry.
Table 4 shows how Mesh-Based Novel View Synthesis allows us to compare against other state-of-the-art methods. We use the differentiable rasterizer from RaDe-GS for training Ours (base). Our approach outperforms 2DGS (Huang et al., 2024) and RaDe-GS (Zhang et al., 2024) across all metrics. GOF (Yu et al., 2024) produces highly detailed meshes, yielding strong PSNR results, but our method achieves comparable PSNR while outperforming in SSIM and LPIPS. This suggests that our meshes are not only accurate but also exhibit fewer visual artifacts and less noise compared to those from GOF. We include a qualitative comparison in the supplementary material.
Moreover, the performance on Tanks & Temples under the Mesh-Based Novel View Synthesis  metric (Table 4) shows strong correlation with the F1 score computed against ground-truth geometry (Table 2). For example, among the common baselines, 2DGS and RaDe-GS rank fourth and third, respectively, under both metrics. Ours (base) and GOF also perform consistently, with the Ours (base) achieving a slightly higher F1 score and GOF performing marginally better on Mesh-Based Novel View Synthesis. These results indicate that our proposed metric is well aligned with ground-truth geometric accuracy.
@@ 6.4. Novel View Synthesis
Figure 7. Effect of adding Normal supervision. Despite showcasing an apparent decrease in performance in our ablation study (see the supplementary material), this figure supports our statement that the combination of depth and normal supervision is essential. The resulting mesh when using only $\mathcal{L}_{\text{MD}}$ is significantly noisier than when supervising with $\mathcal{L}_{\text{mesh}}$.
@@ 6.5. Ablation Study
To evaluate the contribution of each component, we conduct ablation studies on the Tanks and Temples dataset (see Tab. 5). Adding depth supervision via the mesh loss ($\mathcal{L}_{\text{MD}}$) substantially boosts $F_{1}$-score, while further including the normal loss ($\mathcal{L}_{\text{mesh}}$) slightly reduces it. Still, Fig. 7 shows that combining both is critical for removing noise from meshes.
Introducing the anti-erosion loss ($\mathcal{L}_{\text{mesh}}+\mathcal{L}_{\text{erosion}}$) improves reconstruction by preserving thin structures and fine details—especially in challenging regions like fences or vegetation—by penalizing surface erosion. Our loss also acts as a mechanism to recover geometry lost due to SDF sign flips yielding a modest $F_{1}$-score gain (Fig. 5).
@@
Refer to caption Refer to caption
6.6. Analysis of Mesh Extraction
As discussed in Section 2, many existing approaches to mesh extraction from images rely on the TSDF algorithm. These methods typically follow a simple yet seemingly effective pipeline: they first estimate accurate depth maps using a variety of techniques, and then fuse these maps into a mesh via TSDF. However, we argue that TSDF fusion does not scale well to large-scale reconstructions. This limitation arises from its reliance on a fixed 3D grid, which drives up memory requirements as resolution increases, and from its tendency to oversmooth geometric detail.
In Figure 9, we present a quantitative comparison between our extraction method and TSDF fusion. For TSDF fusion, we optimized Gaussians with the same densification strategy as MILo, but removed the mesh-in-the-loop regularization term and employed the same loss function as in (Huang et al., 2024; Zhang et al., 2024; Yu et al., 2024) to avoid bias. TSDF fusion requires substantial memory and is prone to out-of-memory failures; consequently, we exclude the Courthouse scene from this benchmark, as TSDF fusion could not be run across all mesh resolutions due to memory constraints. As shown in Figure 9, our method consistently achieves higher reconstruction quality, demonstrating the robustness of our learnable SDF strategy. We highlight that, TSDF fusion requires iterating over all training viewpoints for each extraction, leading to a long and computationally expensive process that prevents integration into the training loop of Gaussian Splatting representations.
By contrast, MILo leverages learnable SDF values and a scalable pivot set to enable lightweight and efficient mesh extraction, making it practical to incorporate mesh extraction directly into the training loop.
Figure 9. Quantitative evaluation of mesh extraction methods on the Tanks & Temples dataset across different mesh resolutions. We compare (a) traditional TSDF fusion, which relies on a regular 3D grid, with (b) our mesh extraction method, which learns SDF values at Gaussian pivots.
@@ 6.7. Limitations and Conclusions
Future work could explore novel adaptive sampling strategies for Delaunay sites, and improved initialization techniques. Additionally, extending our approach to dynamic scenes and exploring its applications in real-time rendering and interactive applications would be promising directions for future research.
Further exploration of surface-based processing tools that can be integrated into the training pipeline could also yield significant improvements in reconstruction quality and versatility.
@@
Finally, through a a mesh-based novel view synthesis evaluation we highlightM the lack of standardized quantitative protocols to evaluate the alignment of the reconstructed surface with the training images. This is true on widely used datasets such as DTU, MipNeRF 360, and Tanks & Temples. While this approach offers a step toward addressing this gap, it remains an initial effort; future work could focus on refining the evaluation methodology to better capture reconstruction quality in challenging real-world scenes.
7. Acknowledgements
Parts of this work were supported by the ERC Consolidator Grant “VEGA” (No. 101087347), the ERC Advanced Grant “explorer” (No. 101097259), the ANR AI Chair AIGRETTE. Co-funded by the European Union (EU) (ERC Advanced grant FUNGRAPH No 788065 and ERC Advanced Grant NERPHYS No 101141721). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the EU or the European Research Council. Neither the EU nor the granting authority can be held responsible for them.
Supplementary Material
Appendix A Preliminaries
In this supplementary material, we start by briefly reviewing the key concepts that form the foundation of our approach: 3D Gaussian Splatting and Delaunay tetrahedralization.
A.1. 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as a powerful technique for novel view synthesis, offering high-quality rendering with real-time performance. In this representation, a 3D scene is modeled as a collection of 3D Gaussians, each defined by its position $\mu\in\mathbb{R}^{3}$, covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$ (typically parameterized by a scaling vector $s\in\mathbb{R}^{3}$ and a rotation matrix $R\in\mathbb{R}^{3\times 3}$ encoded using a quaternion $q\in\mathbb{R}^{4}$), opacity $\alpha\in(0,1)$, and appearance attributes (such as spherical harmonics coefficients $c\in\mathbb{R}^{S}$ for view-dependent color). During rendering, these 3D Gaussians are projected onto the image plane as 2D Gaussians and composited in a front-to-back order to produce the final image.
A.2. Delaunay Triangulation
Given a set of 3D points, the Delaunay triangulation divides their convex hull into tetrahedra such that no point lies inside the circumsphere of any tetrahedron. This structure has been extensively studied due to its desirable theoretical and practical properties; see (Aurenhammer, 1991) for more details.
In MILo, we rely on the Delaunay triangulation as it naturally adapts to the local density of the input point cloud. Specifically, it creates smaller tetrahedra in regions with higher point density and larger ones in sparser regions. This adaptive nature makes it more efficient than uniform grid-based approaches for extracting surfaces from non-uniform point distributions, such as those formed by optimized 3D Gaussians. Furthermore, the regularity of the Delaunay structure facilitates GPU-accelerated mesh extraction via the marching tetrahedra algorithm (Doi and Koide, 1991), which is the cornerstone of our differentiable mesh extraction process.
Appendix B Experiments
B.1. Implementation Details
Optimization.
For optimization, we use the Adam optimizer with the same learning rates for the Gaussian parameters as (Zhang et al., 2024). We use a learning rate of $0.025$ for the SDF values. We densify Gaussians for 3,000 iterations using aggressive densification (Fang and Wang, 2024b), during which we only rely on our photometric loss. We introduce the volumetric rendering regularization loss ${\mathcal{}L}_{\text{N}}$ after densification, at iteration 3,000. We let Gaussians populate the scene and refine their parameters for 5,000 additional iterations, during which we rely on our volumetric rendering loss ${\mathcal{}L}_{\text{vol}}$. After this point we stop the densification and pruning procedures. Thus, our method does not introduce additional sensitivity to initialization beyond what is already present in standard 3DGS pipelines.
We introduce our mesh extraction pipeline at iteration 8,000. We extract a mesh at every iteration and apply our full loss $\mathcal{}L$} to our representation. The mesh-in-the-loop optimization continues for an additional 10,000 iterations, for a total of 18,000 iterations. For our base model, we prune Gaussians at iteration 8,000 with importance-weighted sampling (Fang and Wang, 2024b) to maintain only the most important Gaussians. For the dense model, we maintain a larger set of Gaussians but compute the importance scores at iteration 8,000 and sample Delaunay sites only from the most important Gaussians. Our weight parameters are $\lambda_{\text{RGB}}=0.2$, $\lambda_{\text{N}}=\lambda_{\text{MD}}=\lambda_{\text{MN}}=0.05$, and $\lambda_{\text{erosion}}=\lambda_{\text{interior}}=0.005$.
Delaunay triangulation.
For the Delaunay triangulation, we use the CGAL library’s 3D Delaunay triangulation implementation, which provides robust and efficient computation of the tetrahedralization. Although Delaunay triangulation is inherently non-differentiable, this does not pose a problem in our setting, as gradients propagate from the mesh back to the Gaussians via the Gaussian pivots. We found that updating the Delaunay triangulation at every iteration is not necessary for stable optimization; We therefore only update the Delaunay triangulation every 500 iterations.
SDF normalization.
In practice, for more stable optimization, we optimize Truncated SDF values normalized to be in the range [1,1][-1,1] using a tanh function. When introducing our Mesh-in-the-Loop optimization, the initial SDF values are computed using a custom, scalable depth-fusion algorithm that operates directly on our Delaunay sites rather than on a regular grid: For each Delaunay site, we compute the signed difference between the depth map rendered from the Gaussians and the depth of the site, and fuse these distances by averaging across all views. This initialization provides a good starting point for the SDF optimization; While being over-smoothed and missing many fine details, the initial mesh is already a reasonable approximation of the surface encoded by the Gaussians.
Mesh rendering.
For differentiable mesh rendering, we use nvdiffrast (Laine et al., 2020), which provides efficient GPU-accelerated rasterization with gradient support. We render the mesh at the same resolution as the Gaussian rendering, using the same camera parameters. To smooth out discontinuities and ensure the propagation of gradients between neighbor pixels, we apply antialiasing smoothing to the depth maps rendered from the mesh.
B.2. Mesh-Based Novel View Synthesis
Figure 10. Our mesh-based evaluation serves as a proxy for quantitatively assessing geometric artifacts, mesh completeness, and background reconstruction. MILo demonstrates superior performance in capturing fine details—such as the intricate chandelier stem and chair spindles—and consistently recovers high-quality background geometry.
We showcase a qualitative comparison of our mesh-based novel view synthesis evaluation in Fig. 10.
B.3. Novel View Synthesis
Table 6. Quantitative results for novel view synthesis on MipNeRF 360 (Barron et al., 2022). We report PSNR, SSIM, and LPIPS. Our method maintains competitive rendering quality while significantly improving surface reconstruction.
Indoor Scenes Outdoor Scenes
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS 30.41 0.920 0.189 24.64 0.731 0.234
Mip-Splatting 30.90 0.921 0.194 24.65 0.729 0.245
BakedSDF 27.06 0.836 0.258 22.47 0.585 0.349
SuGaR 29.43 0.906 0.225 22.93 0.629 0.356
2DGS 30.40 0.916 0.195 24.34 0.717 0.246
GOF 30.79 0.924 0.184 24.82 0.750 0.202
RaDe-GS 30.74 0.928 0.165 25.17 0.764 0.199
Ours (base) 29.96 0.920 0.191 24.47 0.718 0.290
Ours (dense) 30.76 0.934 0.155 24.81 0.744 0.229
While our primary focus is on surface reconstruction, we also evaluate the novel view synthesis quality of our optimized Gaussians. Table 6 presents the PSNR, SSIM, and LPIPS metrics on the MipNeRF 360 (Barron et al., 2022) dataset. Our method maintains competitive rendering quality compared to previous approaches, demonstrating that our mesh-in-the-loop optimization does not compromise the visual fidelity of the Gaussian representation.